OpenStudy (anonymous):

can someone explain how i factor this x^2+10x+16=0

OpenStudy (mimi_x3):

(x+2)(x+8)

OpenStudy (anonymous):

What number times another number equals 16 and adds up to 10? 8 & 2. So you split up the x^2 and get (x+8)(x+2)!

OpenStudy (mimi_x3):

a+b = 2+8 = 10 a*b = 2*8 = 16 xD

OpenStudy (anonymous):

^exactly.(: